OpenStudy (anonymous):

The height of a box with a square base is 8cm and the total surface area of the box is 264cm squared. Find the length of a side of the base of the box

OpenStudy (anonymous):

help pweese

OpenStudy (javk):

You just said that the box has a square base, therefore 'the length of a side of the base of the box' would also be 8cm. If however you were referring to the height of the box, then Step 1: Give the height a variable name let height = h Step 2: Write down the formula used to figure out the surface area of a regular cuboid \[2\times base ^{2}+4\times (base \times height)\] Step 3: Enter all the known values \[2\times 8 ^{2}+4\times (8 times h)=264\] Step 4: Solve for h